# Deep Learning Framework Testing via Heuristic Guidance Based on Multiple Model Measurements

### v1

@@ Abstract
Deep learning frameworks serve as the foundation for developing and deploying deep learning applications. To enhance the quality of deep learning frameworks, researchers have proposed numerous testing methods using deep learning models as test inputs. However, existing methods predominantly measure model bug detection effectiveness as heuristic indicators, presenting three critical limitations: Firstly, existing methods fail to quantitatively measure model’s operator combination variety, potentially missing critical operator combinations that could trigger framework bugs. Secondly, existing methods neglect measuring model execution time, resulting in the omission of numerous models potential for detecting more framework bugs within limited testing time. Thirdly, existing methods overlook correlation between different model measurements, relying simply on single-indicator heuristic guidance without considering their trade-offs. To overcome these limitations, we propose DLMMM, the first deep learning framework testing method to include multiple model measurements into heuristic guidance and fuse these measurements to achieve their trade-off. DLMMM firstly quantitatively measures model’s bug detection performance, operator combination variety, and model execution time. After that, DLMMM fuses the above measurements based on their correlation to achieve their trade-off. To further enhance testing effectiveness, DLMMM designs multi-level heuristic guidance for test input model generation. We apply DLMMM to test three widely used deep learning frameworks (including TensorFlow, PyTorch, and MindSpore). The experimental results show that DLMMM outperforms state-of-the-art methods in effectiveness and efficiency. The test input models generated by DLMMM achieve a 250.13% improvement in operator combination variety. The average model execution time of DLMMM is within 35% of all state-of-the-art methods. DLMMM detects 19 new crashes and 175 new NaN & inconsistency bugs. Among the detected crashes, 16 have been confirmed by developers and two have been fixed in the following version.
Index Terms:
Deep learning framework, framework testing, deep learning model
1 Introduction
@@
Deep Learning (DL) has been widely applied in many safety-critical tasks, such as autonomous driving [1, 2, 3], medical diagnosis [4, 5], and industrial safety monitoring [6, 7]. In recent years, researchers have shown increasing concerns about the security of DL applications [8, 9]. DL frameworks are the foundation for developing and deploying DL applications. Bugs in DL frameworks may potentially lead to application defects [10]. Therefore, the quality of DL frameworks is crucial for the quality of DL applications. To ensure the quality of DL frameworks, researchers propose many framework testing methods. Among them, most generate DL models as test inputs [11, 12]. These methods execute generated models to invoke APIs provided by DL frameworks. During model execution, these methods measure each model’s bug detection effectiveness and heuristically guide the following testing round to promote this measurement. These methods successfully detect many crashes and NaN & inconsistency bugs in DL frameworks. However, there are still three critical limitations in the model measurement design.
Firstly, model’s operator combination variety is not quantitatively measured. Model’s operator combination variety describes the number of substructures contained in a model. Since many framework bugs are triggered by specific model substructures, higher operator combination variety contributes to triggering more unique framework bugs. The above significance of operator combination variety has been recognized by some existing methods (e.g., COMET [13]). However, many DL models contain branches (e.g., ResNet [14]), a special model substructure that enables multi-directional data flow within the model. These models’ operator combination variety cannot be counted properly by existing methods.
Secondly, the model execution time is overlooked. The longer a model’s execution time, the fewer test input models are executed within the same time. Some DL models that might trigger framework bugs may not have the opportunity to be executed, thereby limiting the effectiveness of DL framework testing.
@@
Thirdly, existing methods overlook the correlation between multiple model measurements, focusing narrowly on optimizing one single measurement without achieving overall optimization. For example, the bug detection effectiveness of a single model is a widely used measurement. It measures a model’s effectiveness in triggering framework crashes and NaN & inconsistency bugs. To maximize bug detection effectiveness, existing methods tend to excessively stack more operators into a model. However, this increases the execution time of a model, thereby limiting the efficiency and effectiveness of the overall testing method.
@@
From Finding 1 and 3, it can be seen that higher operator combination variety and shorter model execution time are beneficial for DL framework testing. From Finding 2, it can be seen that is a trade-off is required between increasing operator combination variety and reducing model execution time. Motivated by the above findings, we propose DLMMM, the first DL framework testing method that includes multiple model measurements into heuristic guidance. Firstly, DLMMM quantitatively measures the operator combination variety and model execution time, and adopts the widely used measurement for bug detection effectiveness. Secondly, DLMMM designs a measurement fusion based on the correlation between multiple model measurements to achieve their trade-offs. Thirdly, DLMMM designs a multi-level heuristic, mutation-based test input model generation based on the fused model measurements. In our model generation, the selection of the seed models and the model mutation rules, and the specific implementation of these mutation rules are all heuristically guided.
We conduct experiments to evaluate the effectiveness and efficiency of DLMMM against four state-of-the-art methods, including LEMON [11], Muffin [12], COMET [13] and Gandalf [15]. The experimental results show that DLMMM is better than all state-of-the-art methods in effectiveness and efficiency. Specifically, the average operator combination variety increases by 250.13%. The average model execution time of DLMMM is within 35% of all baselines. DLMMM newly detects 19 crashes and 175 NaN & inconsistency bugs. Among detected crashes, 16 are confirmed by developers, and two are fixed.
@@
  • We implement our method as an open-sourced tool, DLMMM, and release our data on Github.
  • We conduct an evaluation on DLMMM to test three popular DL frameworks. The results show that our method newly detects 19 crashes and 175 NaN & inconsistency bugs, which is the most among all baselines.
More information about DLMMM is available on Github: https://github.com/DeeplearningMMM/DLMMM.
2 Background
2.1 Heuristic DL Framework Testing
@@
Heuristic guidance is an important component in DL framework testing, which is usually used to dynamically control test input model generation. Many DL framework testing methods design heuristic guidance to enhance bug detection effectiveness. For example, COMET [13] applies MCMC [16] search algorithm to generate models covering more operator types. GraphFuzz [17] proposes an operator-level coverage criterion to increase the operator variety. Gandalf [15] applies context-free grammar and deep-Q network to generate test input models with higher model diversity. However, existing methods only focus on the variety of operator types, but fail to generate models with various operator combinations. In addition, the execution time of models generated by existing methods is usually too long. The above challenges motivate us to quantitatively measure operator combination variety and model execution time, and include these measurements into heuristic guidance.
@@
2.2 Representation for Test Input Models in DL Framework Testing
Proper model representation is a prerequisite for model generation. Because graphs clearly represent the relationship between operators and the data flow within a model, Directed Acyclic Graphs (DAG) are widely used to represent test input models in DL framework testing (e.g., LEMON [11] and Muffin [12]). Specifically, each DL model is represented as a DAG, corresponding to a tuple $<V_{G},E_{G}>$. $V_{G}$ denotes the vertex set in the DAG, with each vertex in $V_{G}$ corresponding to a tensor. $E_{G}$ represents the edge set in the DAG, where each edge is a labeled directed edge corresponding to an operator in the model. Specifically, the label of each edge in $E_{G}$ represents the operator type, the starting vertex of the edge represents the input tensor of the operator, and the ending vertex represents the output tensor of the operator. Since the test input model contains only one input tensor and one output tensor, the DAG should include exactly one source vertex and one sink vertex. Furthermore, to ensure the validity of the model, the DAG should be a connected graph.
@@ 2.2 Representation for Test Input Models in DL Framework Testing
3 Empirical Study
We conduct an empirical study to investigate the effect of operator combination variety and model execution time, and the correlation between the two model measurements.
3.1 Empirical Study Setup
We execute four state-of-the-art DL framework testing methods, including LEMON [11], Muffin [12], COMET [13], and Gandalf [15]. Firstly, we download their source code from the open-source repositories. Secondly, we run each method for six hours and save all generated models. Thirdly, we execute saved models under the same environment to test DL frameworks, which include TensorFlow, PyTorch, and MindSpore. For fairness, all test input tensors are randomly generated and kept the same. During model execution, we record each model’s bug detection effectiveness, operator combination variety, and model execution time (see Section 3.2 for their calculation).
Based on recorded model measurements, we investigate their correlation. Considering all recorded measurements are continuous, we use the Pearson correlation coefficient [18] to describe their correlation. The Pearson correlation coefficient is widely used for evaluating the correlation between two variables, with a range of [-1, 1]. The calculation for Pearson correlation coefficient $r$ is:
$r=\frac{\sum_{i=1}^{n}{(X_{i}-\overline{X})(Y_{i}-\overline{Y})}}{\sqrt{\sum_{i=1}^{n}{(X_{i}-\overline{X})^{2}}\sum_{i=1}^{n}{(Y_{i}-\overline{Y})^{2}}}}$
where $n$ represents the number of data points. $X_{i}$ and $Y_{i}$ represent the observed values of the two variables, respectively. $\overline{X}$ and $\overline{Y}$ are the means of the two variables. A positive Pearson correlation coefficient indicates a positive correlation between the two variables, while a negative coefficient indicates a negative correlation. The larger the absolute value of the Pearson correlation coefficient, the stronger the correlation between the two measurements. Generally, when the absolute value of the Pearson correlation coefficient is greater than 0.2, a relatively strong correlation between the two variables is considered to exist.
3.2 Model Measurements
@@
3.2.1 Bug Detection Effectiveness Measurement
@@
Bug detection effectiveness is a widely used measurement by many DL framework testing methods (e.g, Muffin [12], Gandalf [15]). Following these methods, we continue to use this measurement to evaluate the framework testing effectiveness of a single model. The calculation for bug detection effectiveness differs depending on whether the model triggers a crash or NaN bug. Specifically, when the model successfully triggers a crash or NaN bug, the bug detection effectiveness (denoted as $performance$) is the average value of all the data in the test input tensor. If the model fails to trigger a crash or NaN bug, the bug detection effectiveness takes the value of $inconsistency$. $Inconsistency$ is proposed based on the idea of differential testing, which is calculated based on the difference of output tensors under multiple different frameworks. The $inconsistency$ between DL framework $i$ and DL framework $j$ (denoted as $inconsistency_{ij}$) is recorded and calculated as follows. ➀ Provide the same input tensor, perform inference under DL framework $i$ and $j$, and record their result tensor as $result_{i}$ and $result_{j}$. ➁ Calculate the difference between $result_{i}$ and $result_{j}$ as $result\_diff_{ij}$. ➂ Calculate the maximum value among all numbers in $result\_diff_{ij}$ as $inconsistency_{ij}$.
3.2.2 Operator Combination Variety Measurement
@@
Operator-level variety (including operator variety and operator combination variety) has attracted widespread attention from DL framework testing researchers. Many researchers (e.g., Gandalf [15], Muffin [12]) qualitatively believe that since many DL framework bugs are closely related to specific model substructures, the higher the variety of operators and operator combinations in the model, the more effective the framework testing will be. However, although operator combination variety is crucial for enhancing the effectiveness of framework testing, designing a universal quantitative measurement for operator combination variety is a highly challenging task. The main challenge lies in the complexity of model structures, as some models include branches. Existing methods have made some attempts but have failed to overcome this challenge. For example, COMET [13] measures operator combination variety based on the operator sequence of the model. Operator sequence can only describe single-directional data flow within the model, instead of representing model with branches (multi-directional data flow). A more universal quantitative measurement is urgently needed.
@@
To overcome this challenge, we propose the measurement $variety\_degree$ for operator combination variety. $Variety\_degree$ is calculated based on the number of subgraphs contained in the DAG of the model. The detailed definition and calculation for $variety\_degree$ is as follows: Firstly, set the number of operators allowed in a subgraph, called $depth$. Secondly, find all subgraphs, called $motifs$, in the DAG representation of the model, where the number of operators in each subgraph equals to $depth$. Thirdly, remove all isomorphic subgraphs from $motifs$. Finally, count the remaining subgraphs in $motifs$, called $variety\_degree$.
@@
$variety\_degree$ is a universal measurement that reflects the variety of both single operators and operator combinations. Specifically, when $depth=1$, $variety\_degree$ describes the variety of single operators in the model; And when $depth>1$, $variety\_degree$ describes the variety of operator combinations in the model. Since operator combinations with fewer operators often have higher practical reproducible value, the $depth$ is set to 2.
3.2.3 Model Execution Time Measurement

### v2

@@ Abstract.
Deep learning frameworks serve as the foundation for developing and deploying deep learning applications. To enhance the quality of deep learning frameworks, researchers have proposed numerous testing methods using deep learning models as test inputs. However, existing methods predominantly measure model bug detection effectiveness as heuristic indicators, presenting three critical limitations. Firstly, existing methods fail to quantitatively measure model’s operator combination variety, potentially missing critical operator combinations that could trigger framework bugs. Secondly, existing methods neglect measuring and heuristically guiding the model execution time, resulting in the omission of numerous models potential for detecting more framework bugs within limited testing time. Thirdly, existing methods overlook correlation between different model measurements, relying simply on single-indicator heuristic guidance without considering their trade-offs. To overcome these limitations, we propose DLMMM, the first deep learning framework testing method to include multiple model measurements into heuristic guidance and fuse these measurements to achieve their trade-offs. DLMMM firstly quantitatively measures model’s bug detection performance, operator combination variety, and model execution time. After that, DLMMM fuses these measurements based on their correlation to achieve their trade-offs. To further enhance testing effectiveness, DLMMM designs multi-level heuristic guidance for test input model generation. We apply DLMMM to test three widely used deep learning frameworks (including TensorFlow, PyTorch, and MindSpore). The experimental results show that DLMMM outperforms state-of-the-art methods in effectiveness and efficiency. The test input models generated by DLMMM achieve a 250.13% improvement in operator combination variety. The average model execution time of DLMMM is within 35% of all state-of-the-art methods. DLMMM detects 19 new crashes and 175 new NaN & inconsistency bugs. Among detected crashes, 16 have been confirmed by developers and two have been fixed in the following version.
Deep Learning, Framework Testing, Heuristic Guidance
copyright: acmcopyrightjournalyear: 2025isbn: XXXXX-XXXX-XXXXX-XXXX-Xccs: Software Engineering Software Testing
1. Introduction
@@
Deep Learning (DL) has been widely applied in many safety-critical tasks, such as autonomous driving (autodrive, ; autodrive2, ; autodrive3, ), medical diagnosis (disease_diagnosis_1, ; disease_diagnosis_2, ), and industrial safety monitoring (industrialsafetymonitoring1, ; industrialsafetymonitoring2, ). In recent years, researchers have shown increasing concerns about the security of DL applications (DLSecurity1, ; DLSecurity2, ). DL frameworks are the foundation for developing and deploying DL applications. Bugs in DL frameworks may potentially lead to application defects (relation_DLapplication_framework, ). Therefore, the quality of DL frameworks is crucial for the quality of DL applications. To ensure the quality of DL frameworks, researchers propose many framework testing methods. Among them, most generate DL models as test inputs (lemon, ; muffin, ). These methods execute generated models to invoke APIs provided by DL frameworks. To generate more effective models, these methods design measurements for each model’s bug detection effectiveness and heuristically guide the following testing round to promote this measurement. These methods successfully detect many crashes and NaN & inconsistency bugs in DL frameworks. However, there are still three critical limitations in the model measurement design.
Firstly, the diversity of operator combinations within a test input model lacks the quantitative measurement. This diversity refers to the number of distinct substructures in a model, where each substructure represents a specific way of DL framework invocation. A greater variety of operator combinations implies more diverse ways in which the framework is invoked, ultimately leading to more comprehensive testing. Due to the complex structure of the model, quantitatively measuring the operator combinations variety is a significant challenge. Specifically, many DL models contain branches (e.g., ResNet (resnet, )), a special model substructure that enables multi-directional data flow within the model. Such models with branches cannot be simply represented as operator sequences, but require representation via directed acyclic graphs (DAGs). Accordingly, operator sequence-based variety measurements are no longer applicable to these models. To assess the variety of such structurally complex models, a new DAG-based measurement method needs to be designed.
Secondly, since the model execution time is not quantitatively measured or incorporated into heuristic guidance, the average model execution time is too long, which limits testing effectiveness. This limitation arises because a single model execution can trigger at most one DL framework bug—once triggered, this bug’s impact propagates to all subsequent operators, preventing the exposure of other bugs during this execution. A longer average model execution time reduces the number of test input models that can be executed within the same testing time. As a result, some potentially bug-revealing models do not have the opportunity to be executed within the limited testing time, ultimately leading to some bugs not being successfully triggered.
@@
Thirdly, existing methods overlook the correlation between multiple model measurements, focusing narrowly on optimizing one single measurement without achieving overall optimization. For example, the bug detection effectiveness of a single model is a widely used measurement. It measures a model’s effectiveness in triggering framework crashes and NaN & inconsistency bugs. To maximize bug detection effectiveness, existing methods tend to excessively stack more operators into a model. However, this increases another important measurement, the average model execution time, thereby limiting the efficiency and effectiveness of the overall testing method.
@@
From Finding 1 and 3, higher operator combination variety and shorter model execution time are beneficial for DL framework testing. From Finding 2, a trade-off is required between increasing operator combination variety and reducing model execution time. Motivated by the above findings, we propose DLMMM, the first DL framework testing method that includes multiple model measurements into heuristic guidance. Firstly, DLMMM quantitatively measures the operator combination variety and model execution time, and adopts the widely used measurement for bug detection effectiveness. Secondly, DLMMM designs a measurement fusion based on the correlation between multiple model measurements to achieve their trade-offs. Thirdly, DLMMM designs a multi-level heuristic, mutation-based test input model generation based on the fused model measurements. In our model generation, the selection of the seed models and the model mutation rules are all heuristically guided.
We conduct experiments to evaluate the effectiveness and efficiency of DLMMM against four state-of-the-art methods, including LEMON (lemon, ), Muffin (muffin, ), COMET (comet, ) and Gandalf (gandalf, ). The experimental results show that DLMMM is better than all state-of-the-art methods in effectiveness and efficiency. Specifically, the average operator combination variety increases by 250.13%. The average model execution time of DLMMM is within 35% of all baselines. DLMMM newly detects 19 crashes and 175 NaN & inconsistency bugs. Among detected crashes, 16 are confirmed by developers, and two are fixed.
@@
  • We implement our method as an open-sourced tool, named DLMMM, and conduct an evaluation on DLMMM to test three popular DL frameworks. The results show that our method newly detects 19 crashes and 175 NaN & inconsistency bugs, which is the most among all baselines. In addition, DLMMM enhances operator combination variety by 250.13% over the best baseline while maintaining an average model execution time within 35% of state-of-the-art methods.
More information is available on Github: https://github.com/DeepLMMM/DLMMM.
2. Background
2.1. Heuristic DL Framework Testing
@@
Heuristic guidance is an important component in DL framework testing, which is usually used to dynamically control the test input model generation. Many DL framework testing methods design heuristic guidance to enhance bug detection effectiveness. For example, COMET (comet, ) applies MCMC (MCMC, ) search algorithm to generate models covering more operator types. GraphFuzz (fanggraphbased, ) proposes an operator-level coverage to increase the operator variety. Gandalf (gandalf, ) applies context-free grammar and deep-Q network to generate test input models with higher model diversity. However, existing methods only focus on the variety of operator types, but fail to generate models with various operator combinations. In addition, the execution time of models generated by existing methods is usually too long. The above challenges motivate us to quantitatively measure operator combination variety and model execution time, and include these measurements into heuristic guidance.
@@
2.2. Representation for Test Input Models in DL Framework Testing
Proper model representation is a prerequisite for model generation. Because graphs clearly represent the relationship between operators and the data flow within a model, Directed Acyclic Graphs (DAG) are widely used to represent test input models in DL framework testing (e.g., LEMON (lemon, ) and Muffin (muffin, )). Specifically, each DL model is represented as a DAG, corresponding to a tuple $V_{G},E_{G}>$. $V_{G}$ denotes the vertex set in the DAG, with each vertex in $V_{G}$ corresponding to a tensor. $E_{G}$ represents the edge set in the DAG, where each edge is a labeled directed edge corresponding to an operator in the model. Specifically, the label of each edge in $E_{G}$ represents the operator type, the starting vertex of the edge represents the input tensor of the operator, and the ending vertex represents the output tensor of the operator. Since the test input model contains only one input tensor and one output tensor, the DAG should include exactly one source vertex and one sink vertex. Furthermore, to ensure the validity of the model, the DAG should be a connected graph.
@@ 2.2. Representation for Test Input Models in DL Framework Testing
3. Empirical Study
As mentioned in Section 1, researchers’ subjective experience suggests that operator combination variety and model execution time are potentially important measurements that may influence bug detection performance. However, the impact of these measurements has not been quantitatively characterized. Moreover, existing studies have not explored the trade-offs between these measurements. To bridge this gap and lay the groundwork for designing DLMMM’s heuristic guidance, we conduct an empirical study to investigate the effects of operator combination variety and model execution time, as well as the correlation between these model measurements.
3.1. Empirical Study Setup
We execute four state-of-the-art DL framework testing methods, including LEMON (lemon, ), Muffin (muffin, ), COMET (comet, ), and Gandalf (gandalf, ). Firstly, we download their source code from the open-source repositories. Secondly, we run each method for six hours and save all generated models. No new bugs are detected in the last hour, indicating the convergence of the experiment. Thirdly, we execute saved models under the same environment to test DL frameworks, which include TensorFlow, PyTorch, and MindSpore. For fairness, all test input tensors are randomly generated and kept the same. During the model execution, we record each model’s bug detection effectiveness, operator combination variety, and model execution time (see Section 3.2 for their calculation).
Based on recorded model measurements, we investigate their correlation. Considering the potential non-normal distribution of the data and the presence of outliers, we use the Spearman correlation coefficient to describe their relationship. The Spearman correlation coefficient is a measurement used for evaluating the correlation between two variables, with a range of [-1, 1] (Spearman_correlation_coefficient, ). The calculation for the Spearman correlation coefficient $\rho$ is:
$\rho=1-\frac{6\sum_{i=1}^{n}d_{i}^{2}}{n(n^{2}-1)}$
where $n$ represents the number of data points, and $d_{i}$ represents the difference between the ranks of corresponding values of the two variables. A positive Spearman correlation coefficient indicates a positive monotonic correlation between the two variables, while a negative coefficient indicates a negative monotonic correlation. The larger the absolute value of the Spearman correlation coefficient, the stronger the monotonic correlation between the two measurements. Generally, when the absolute value of the Spearman correlation coefficient is greater than 0.2, a relatively strong monotonic correlation between the two variables is considered to exist (Spearman_correlation_coefficient, ).
3.2. Model Measurements
@@
3.2.1. Bug Detection Effectiveness Measurement
@@
Bug detection effectiveness is a widely used measurement by many DL framework testing methods (e.g, Muffin (muffin, ), Gandalf (gandalf, )). Following these methods, we continue to use this measurement to evaluate the framework testing effectiveness of a single model. The calculation for bug detection effectiveness differs depending on whether the model triggers a crash or NaN bug. Specifically, when the model successfully triggers a crash or NaN bug, the bug detection effectiveness (denoted as performanceperformance) is the average value of all the data in the test input tensor. If the model fails to trigger a crash or NaN bug, the bug detection effectiveness takes the value of inconsistencyinconsistency. InconsistencyInconsistency is proposed based on the idea of differential testing, which is calculated based on the difference of output tensors under multiple different frameworks. The inconsistencyinconsistency between DL framework $i$ and DL framework $j$ (denoted as $inconsistency_{ij}$) is recorded and calculated as follows. ➀ Provide the same input tensor, perform inference under DL framework $i$ and $j$, and record their result tensor as $result_{i}$ and $result_{j}$. ➁ Calculate the difference between $result_{i}$ and $result_{j}$ as $result\_diff_{ij}$. ➂ Calculate the maximum value among all numbers in $result\_diff_{ij}$ as $inconsistency_{ij}$.
3.2.2. Operator Combination Variety Measurement
@@
Operator-level variety (including operator variety and operator combination variety) has attracted widespread attention from DL framework testing researchers. Many researchers (e.g., Gandalf (gandalf, ), Muffin (muffin, )) qualitatively believe that since many DL framework bugs are closely related to specific model substructures, the higher the variety of operators and operator combinations in the model, the more effective the framework testing will be. However, although operator combination variety is crucial for enhancing the effectiveness of framework testing, designing a universal quantitative measurement for operator combination variety is a highly challenging task. The main challenge lies in the complexity of model structures, as some models include branches. Existing methods have made some attempts but have failed to overcome this challenge. For example, COMET (comet, ) measures operator combination variety based on the operator sequence of the model. Operator sequence can only describe single-directional data flow within the model, instead of representing model with branches (multi-directional data flow). A more universal quantitative measurement is urgently needed.
@@
To overcome this challenge, we propose the measurement variety_degreevariety\_degree for operator combination variety. Variety_degreeVariety\_degree is calculated based on the number of subgraphs contained in the DAG of the model. The detailed definition and calculation for variety_degreevariety\_degree is as follows: Firstly, set the number of operators allowed in a subgraph, called depthdepth. Secondly, find all subgraphs, called motifsmotifs, in the DAG representation of the model, where the number of operators in each subgraph equals to depthdepth. Thirdly, remove all isomorphic subgraphs from motifsmotifs, which are considered as redundant subgraphs. Finally, count the remaining subgraphs in motifsmotifs, called variety_degreevariety\_degree.
@@
variety_degreevariety\_degree is a universal measurement that reflects the variety of both single operators and operator combinations. Specifically, when $depth=1$, variety_degreevariety\_degree describes the variety of single operators in the model; And when $depth>1$, variety_degreevariety\_degree describes the variety of operator combinations in the model. Since operator combinations with fewer operators often have higher practical reproducible value, and it is too time-consuming to set the depthdepth larger than 2, the depthdepth is set to 2.
3.2.3. Model Execution Time Measurement